How do I do this:
Solve the equation 2/5 - 2/15x= 14/15
I can't figure that out no matter how hard I try HELP!!!

Answer by checkley75(3666)

2/5-2/15x=14/15
-2/15x=14/15-2/5
-2/15x=(14-3*2)/15
-2/15x=(14-6)/15
-2x=8
x=8/-2
x=-4 answer.
2/5-(2/15)-4=14/15
2/5+8/15=14/15
(2*3+8)/15=14/15
(6+8)/15=14/15
14/15=14/15